OpenStudy (anonymous):

Michele wanted to measure the height of her school’s flagpole. She placed a mirror on the ground 48 feet from the flagpole, then walked backward until she was able to see the top of the pole in the mirror. Her eyes were 5 feet above the ground and she was 12 feet from the mirror. Using similar triangles, find the height of the flagpole to the nearest tenth of a foot.

OpenStudy (anonymous):

Let the height of the flagpost be h. Then, h/48 = 5/12 h = 20 feet.

OpenStudy (anonymous):

ya u looked tht up ^

OpenStudy (anonymous):

ya lul

OpenStudy (anonymous):

lmao i did same XD didnt feel like reading question

OpenStudy (anonymous):

im just tryna get a medal

OpenStudy (anonymous):

it is 20 tho

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ya my other account just got banned for showing work but giving answers so watch out

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i had a 62 score :( and im bored now cuz im suppose to be doing work lol its banned for 2 days

OpenStudy (anonymous):

i fanned you @2_muh_$w3G

OpenStudy (anonymous):

nooo dont turn this into a chat XD

OpenStudy (anonymous):

ok hey can i has a medal

OpenStudy (anonymous):

oh and thanx for ur help

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

how do u get medals or send them?

OpenStudy (anonymous):

you click on best response in a question

OpenStudy (anonymous):

oh that sound simple enough...i did it :D

OpenStudy (anonymous):

nice